OpenStudy (anonymous):

The new business building uptown is building a wall made entirely of glass pyramids. Each pyramid will have a square base with length of 8 inches and height of 10 inches. If the wall requires 250 of these pyramids, what volume of glass will be used? Round to the nearest cubic inch.

OpenStudy (anonymous):

If a pyramid has a base of B square units, and a height of h units, then the a Pyramid volume V is 1/3 B x h cubic units, or V = 1/3 Bh ... the Base has to be AREA so with a square base you would need to square the side length to be in squared form for this it needs to be V=1/3(8)^2 X (10) = 1/3(64)(10) ............ = 21.12 x 10 = ..........211cm^3

OpenStudy (anonymous):

is that rounded ..

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

I cant put that in my test it wouldnt take the ^ sign

OpenStudy (anonymous):

is there another way to write it

OpenStudy (anonymous):

@shannon342

OpenStudy (anonymous):

are you taking the test online or its hand written

OpenStudy (anonymous):

i tried just 211 and it counted it wrong & online

OpenStudy (anonymous):

\[211cm ^3\]

OpenStudy (anonymous):

are u able to write it like this

OpenStudy (anonymous):

are u can put it in words

OpenStudy (anonymous):

one sec and you dont need the unit

OpenStudy (anonymous):

211cm cubed

OpenStudy (anonymous):

and no it has to be numbers

OpenStudy (anonymous):

did it work ?

OpenStudy (anonymous):

no /:

OpenStudy (anonymous):

it should take ^ sign idk why it isnt taking it

OpenStudy (anonymous):

idk why maybe your answer is wrong

OpenStudy (anonymous):

Nope answer is right

OpenStudy (anonymous):

why is it cm ?

OpenStudy (anonymous):

and u sure you rounded correctly

OpenStudy (anonymous):

Centimeter

OpenStudy (anonymous):

well we dont need that cuz on all the other problems i still got them right with no unit